OpenStudy (anonymous):

HELP pleaseeeeee :)) Explain, in complete sentences, how to factor 4x2 – 4x – 48. Explain how your factors could be checked for accuracy.

OpenStudy (mertsj):

Four is a common factor. Factor it out. The remaining trinomial is the product of two binomials. Factor it. Check your answer by multiplication.

OpenStudy (anonymous):

4(x^2 - x - 12) = 0 4(x + 3)(x - 4) = 0 => multiply to check

OpenStudy (mertsj):

I guess this question has been answered already.

OpenStudy (anonymous):

i dont understand it :/

OpenStudy (mertsj):

Four is a common factor. Do you understand that?

OpenStudy (anonymous):

factor 4x^2 – 4x – 48. factor out 4 in each of the three terms 4(x^2 - x - 12) Find the factors of 12 which differ by 1, -x means the larger factor will be neg 4(x-4)(x+3) : You can confirm this by FOILing (x-4)(x+3) and multiplying that by 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can divide the equation by 4 BUT you don't want to change the equation by 4, sooo what do you do? you also multiply it by 4 therefore you get, \[4(x^2-x-12)=y\] now you need to find two numbers that multiply together to get you -12 but also when you add them together you get -1 so notice that (-4)*(3)=-12 and 3-4=-1 so you can do whats called factor by grouping --> \[y=4x^2-4x-48\\=4(x^2-x-12)\\=4(x^2-4x+3x-12)\\=4((x^2-4x)+(3x-12))\\=4(x(x-4)+3(x-4))\\=4(x+3)(x-4)\]

OpenStudy (mertsj):

So factor out the 4

OpenStudy (anonymous):

oh geeze it took me so long to type that that I didn't realise the q had already been answered lol

OpenStudy (anonymous):

do u understand now?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yes. i do thank you! :)